OpenStudy (anonymous):

The mean of seven numbers is 29. If one number is discarded the mean is now 33. What is the number that has been discarded?

OpenStudy (anonymous):

mean of 7 numbers=29 total of 7 numbers=29*7=203 mean of 6 numbers=33 total of 6 numbers=33*6=198 discarded number=203-198=?